OpenStudy (anonymous):

Use division to determine whether the ratios form a proportion. 1/4 6/24 and A. Yes. They form a proportion. B. No. They do not form a proportion.

OpenStudy (anonymous):

what do I divide like do i do 1/6? and 4/24?

OpenStudy (anonymous):

1/6=6. and 24/4=6 6/6=1? Like that

OpenStudy (anonymous):

@Data_LG2

OpenStudy (anonymous):